OpenStudy (anonymous):

HHHHHHEEEEELLLLLLLPPPPPPP cos^-1(1)........ i know the answer is 0 but why?

OpenStudy (anonymous):

because \(\cos(0)=1\) and it is the only number between \(0\) and \(\pi\) for which that is true

OpenStudy (anonymous):

is there a better explanation?

OpenStudy (anonymous):

not really, no

OpenStudy (anonymous):

you are looking for a number \(x\) in the interval \([0,\pi]\)with \(\cos(x)=1\) that number is \(0\) and that is all there is to it

myininaya (myininaya):

that would be my best explanation too